OpenStudy (anonymous):

If x-1, x, x-1 is a geometric sequence, then x= (a) -1/2 (b) 1/2 (c) 2 (d) 1 (e) -1

OpenStudy (anonymous):

c i think. might have someone else check that

OpenStudy (anonymous):

just to be shure

OpenStudy (anonymous):

how'd u get the answer?

OpenStudy (anonymous):

oh wait thats wrong its -1

OpenStudy (anonymous):

because x by itself is 1. so x=-1 +x would =0 then add the last -1 that =-1

OpenStudy (anonymous):

geometric sequence means the ratio of any 2 consecutive terms is the same. so (x-1/x)=(x/x-1) now solve..

OpenStudy (anonymous):

i got x = 1/2

OpenStudy (anonymous):

yes right

OpenStudy (anonymous):

sorry i was wrong

OpenStudy (anonymous):

it's ok and thanks to the both of you